# Grand-Canonical Symmetric Orbifold Theories

### v1

@@ Abstract
In this letter, we study grand-canonical symmetric orbifolds of conformal field theories on the sphere. We show that a consistent operator product expansion can be defined, provided the Hilbert space is taken to be the direct sum of symmetric orbifolds of all degrees. This Hilbert space contains a tower of central operators of conformal dimension zero, one of which represents the central charge of the Virasoro algebra. Our construction thus provides a conformal field theory interpretation of the central charge operator $\mathcal{I}$ in $AdS_{3}$ string theory.
@@ I Introduction
String theory on an $AdS_{3}\times S^{3}\times T^{4}$ background with one unit of NS-NS flux is dual to a grand-canonical ensemble of symmetric orbifold conformal field theories [1, 2, 3, 4, 5]. So far, these ensembles have been defined solely through their correlation functions. Up to a normalization factor, they read
$\langle\cdot\rangle_{p}:=\sum_{d}p^{d}\langle\cdot\rangle^{un}_{d}\,$, (1)
where $\langle\cdot\rangle^{un}_{d}$ denotes the unnormalized correlation function in the symmetric orbifold of degree $d$. The purpose of this letter is to investigate these ensembles further and show that they realize much of the structure of a single conformal field theory. In particular, we prove that operators in a grand-canonical symmetric orbifold admit an operator product expansion. Its consistency requires the Hilbert space to be enlarged beyond that of a single symmetric orbifold in the large $N$ limit. The Hilbert space transforms in a reducible representation of the Virasoro algebra that is not completely reducible. Its central charge is represented by a non-scalar operator, which we identify with the central charge operator $\mathcal{I}$ in the dual string theory. Our construction therefore sheds light on the nature of the operator $\mathcal{I}$ from a conformal field theory perspective.
@@ I.2 Road map
where $Z_{p}$ denotes the partition function of the grand-canonical theory. At each degree, one can take the operator product expansion of the twist operators. However, the corresponding OPE coefficients depend on the degree. For example, $C_{kk}^{\quad 0}(d)=\frac{d!}{k(d-k)!}$. Therefore, it is unclear whether an operator product expansion can be defined in the grand-canonical theory. To prove that such an expansion would be inconsistent, let us compute the grand-canonical two-point function explicitly:
@@
An alternative way to describe the central charge operator $\mathcal{I}$ is as a differential operator with respect to the chemical potential [5]. As such, it generates an algebra of operators absent from the Hilbert space of the symmetric orbifold. Therefore, if the central charge operator of the string theory admits a field-theoretic dual, then its description requires the Hilbert space of the grand-canonical theory to be substantially enlarged. Following the intuition gained from the study of topological symmetric orbifolds [12, 13, 14], we will propose that the algebraic structure required for this enlargement is that of partial permutations [15]. A partial permutation is a permutation of a subset of $n$ elements. It is composed of a support $\mathcal{S}\subset\{1,\dots,n\}$ and a permutation $\rho$ of $\mathcal{S}$. We denote it as $(\mathcal{S},\rho)$, or alternatively with the Latin letter $r$. The product of two partial permutations is obtained by multiplying the permutations and concatenating their supports:
@@
We define the orbit sum $A_{[r]}$ as the sum over all partial permutations in the conjugation orbit of $r$. These orbit sums are characterized by the size of their support and their cycle type. Our proposal is that in a grand-canonical theory, twist operators are indexed not by orbits of the infinite symmetric group, but rather by orbits of partial permutations. An immediate consequence of this proposal is that there are many central operators. In section III, we will identify these operators within the algebra generated by the dual of the central charge operator $\mathcal{I}$. In the next section, we focus on clarifying the structure of the grand-canonical theory: we build up the Hilbert space, explain how a consistent operator product expansion can be defined, and determine the representation of the Virasoro algebra.
@@ II.1 The Hilbert Space
To construct the Fock space of quantum field theory, one must sum over all $n$-particle Hilbert spaces. Similarly, the Hilbert space of the grand-canonical symmetric orbifold theory is obtained by summing the Hilbert spaces corresponding to all degrees:
$\mathcal{H}=\bigoplus_{d\geq 0}\mathcal{H}_{d}\,.$ (15)
Each Hilbert space is itself the sum of twisted sectors. We may therefore write
@@
where the sum runs over orbits of partial permutations of all sizes. We denote by $\alpha_{[r]}$ the twist operator that generates the ground state of the sector $\mathcal{H}_{[r]}$, and refer to it as a generalized twist operator. The partial permutation orbit of empty support $\alpha_{\emptyset}$ is the identity operator. It corresponds under the state-operator map to the vacuum of $\mathcal{H}_{0}$, which should therefore be regarded as the vacuum of the theory.
It is convenient to organize the Hilbert space in terms of the boundary conditions its fields obey. Given a partial permutation $r$ with minimal support $\mathcal{S}_{r}$, we define
$\hat{\mathcal{H}}_{r}:=\bigoplus_{k\geq 0}\mathcal{H}_{[(\mathcal{S}_{r}\cup 1^{k},r)]}\,.$ (17)
The Hilbert space $\mathcal{H}$ is the direct sum of the $\hat{\mathcal{H}}_{r}$.
@@ II.2 Correlation Functions
We now explain how to compute disconnected correlation functions in the grand-canonical theory. We define a fixed-degree correlation function that involves a generalized twist field as
$\langle\alpha_{(\mathcal{S},\rho)}\cdots\rangle_{d}:=\langle\sigma_{\rho}\cdots\rangle^{un}_{d}\,.$ (18)
Although these correlators match, the corresponding gauge-invariant observables differ by a combinatorial factor. For example,
$\langle\alpha_{k,1}(z)\alpha_{k}(0)\rangle_{d}$ $=\frac{1}{k(d-k-1)!}\cdot\frac{Z_{seed}^{d}}{z^{2h_{k}}}\,$, (19)
which differs from $\langle\sigma_{k}\sigma_{k}\rangle_{d}$ by a factor dkd-k. This factor accounts for the larger number of elements entering the definition of $\alpha_{k,1}$, as compared to $\sigma_{k}$. In general, we find [15]:
$\langle\alpha_{[r]}(z)\cdot(\dots)\rangle_{d}=\binom{m_{1}(\rho)}{m_{1}(r)}\langle\sigma_{[\rho]}(z)\cdot(\dots)\rangle_{d}\,$, (20)
where $m_{1}(r)$ is the number of trivial elements in the support of the partial permutation $r$, and $m_{1}(\rho)$ the number of trivial elements in the corresponding permutation $\rho$ of $\{1,2,\dots,d\}$. Geometrically, the operator $\alpha_{[r]}$ selects $\lvert\mathcal{S}_{r}\rvert$ sheets out of the $d$ available in the covering space. Only those sheets contribute to connected correlators. This geometric picture is perhaps best understood from the dual string perspective, where trivial elements in the support of a partial permutation correspond to disjoint worldsheets in their vacuum state, see section III.4.
Additionally, we note that combinatorial differences in fixed-degree correlation functions can lead to drastically different behaviours once summed over all degrees. For example, the operator $\alpha_{1}$ behaves within any fixed-degree correlation function as a factor of the degree:
@@
On the other hand, the generalized twist operators $\alpha_{k\geq 2}$ behave precisely as the standard $\sigma_{k}$ in all correlation functions, and we may therefore view the $\alpha_{k}$ as the standard twist operators in the grand-canonical theory.
Finally, we remark that the dependence of grand-canonical correlation functions on the positions of the fields always factors out of the sum over degrees, as exemplified by equation (9). It follows that the grand-canonical correlation functions inherit the conformal invariance of the fixed-degree correlators.
@@ II.3 Operator Product Expansion
In order to take operator product expansions in the grand-canonical theory, we first compute at a given degree $d$, then replace each factor of $d$ by the operator $\alpha_{1}$. As we will see in section III.3, the product of $\alpha_{1}$ with any other twist operator can be calculated exactly. Multiplying out its powers leads to the desired expansion, which is manifestly independent of the degree. On the one hand, this immediately implies that the operator product expansion is consistent, because it factorizes out the grand-canonical sum. On the other hand, it implies the existence of a well-defined operator product expansion in the large $N$ limit of the fixed-degree symmetric orbifold. In the next section, we study the self-OPE of the stress-energy tensor of the theory. It provides a concrete example of the technique laid down above.
II.4 Virasoro Algebra and Central Charge
The stress-energy tensor of the grand-canonical theory coincides with the stress-energy tensor of a single symmetric orbifold in the large $N$ limit:
@@ II.4 Virasoro Algebra and Central Charge
where the $L_{n}$ are the Laurent modes of the stress-energy tensor, $L_{n}:=\int dz\,z^{n+1}T(z)$. An infinite-dimensional generalization of Schur’s lemma implies that the representation at hand is reducible, as it admits a non-scalar central element. In fact, the grand-canonical Hilbert space exhibits the structure of a descending filtration of reducible representations:
$\bigoplus_{d\geq 0}\mathcal{H}_{d}\supset\bigoplus_{d\geq 1}\mathcal{H}_{d}\supset\bigoplus_{d\geq 2}\mathcal{H}_{d}\supset\dots$ (26)
Nevertheless, the theory does not decompose into superselection sectors because the representation is not completely reducible. Indeed, the commutation relations of the Virasoro algebra generate the central charge operator, which is of degree one in the filtration. The Hilbert spaces $\hat{\mathcal{H}}_{r}$ defined in (17) are the indecomposable building blocks of the theory.
To conclude, we note that the appearance of a non-scalar central charge in the representation (25) is consistent with the findings of [16] for the spacetime Virasoro algebra of $AdS_{3}$. In the tensionless context, it leads to the identification of $\alpha_{1}$ as the counterpart of the central charge operator $\mathcal{I}$ of the dual string theory. The expectation value of $c_{seed}\,\alpha_{1}$ therefore computes the Brown-Henneaux central charge [17]. This allows us to precisely identify the relation between the string coupling and the chemical potential. The three-dimensional Newton constant for an $AdS_{3}\times S^{3}\times T^{4}$ background at $k=1$ reads [18]
$G_{N}^{(3)}=\frac{G_{N}^{(10)}}{\text{Vol}(S^{3})\cdot\text{Vol}(\mathbb{T}^{4})}=\frac{8\pi^{6}g_{s}^{2}R^{8}}{2\pi^{2}R^{3}\cdot(2\pi)^{4}v_{4}R^{4}}\,$, (27)
@@
II.5 A bona-fide conformal field theory ?
At this stage, let us summarize which conformal field theory axioms the grand-canonical symmetric orbifold satisfies. Firstly, the conformal invariance is manifest. Secondly, the state-operator correspondence and commutativity axioms hold by construction. Thirdly, we have shown that a consistent operator product expansion can be defined. Nevertheless, we do not expect grand-canonical ensembles to be well-defined on Riemann surfaces of genus $g\geq 2$, due to the rapid growth of the handle operator with the degree. In fact, these theories fail to satisfy the sewing axiom of [22]. Moreover, their spectrum features reducible but indecomposable representations. Such representations are emblematic of logarithmic conformal field theories. In these theories, the dilation operator is not diagonalizable, which leads to logarithmic singularities for the correlation functions [23]. The central charge, on the other hand, is represented by a multiple of the identity. Interestingly, the grand-canonical symmetric orbifold satisfies the same axioms as a logarithmic conformal field theory on the sphere, but shares neither of the properties described above. The takeaway is that the axiom of decomposability can be violated in qualitatively distinct ways.
@@ III.1 The Central Charge Operator
$\langle\alpha_{1}\cdot(\dots)\rangle^{un}$ $:=\sum_{d\geq 1}d\,p^{d}\langle(\dots)\rangle^{un}_{d}$ (30)
$=p\partial_{p}\langle(\dots)\rangle^{un}\,.$ (31)
@@ III.2 The Universal Subalgebra
$\langle\alpha_{1^{k}}\cdot(\dots)\rangle^{un}$ $:=\sum_{d\geq 1}\binom{d}{k}p^{d}\langle(\dots)\rangle^{un}_{d}\,$, (33)
$=\frac{p^{k}\partial_{p}^{k}}{k!}\langle(\dots)\rangle^{un}\,.$ (34)
@@
Finally, we propose to view the existence of this algebra as a diagnostic for determining whether a theory is grand-canonical: a conformal field theory is grand-canonical if and only if it contains an algebra freely generated by a single central operator.
@@ III.3 OPE with the Central Charge Operator
with $k\geq 2$. A geometric interpretation of equation (44) is that $\alpha_{1}$ either marks an existing sheet or creates a new one. Holographically, the latter corresponds to adding a disjoint worldsheet in its vacuum state. Indeed, the Hilbert space (16) distinguishes between vacuum configurations with different numbers of worldsheets. A key takeaway of our construction is that for the boundary description to exhibit a field theory structure, it must holographically encode a multi-string Hilbert space.

### v2

@@ Abstract
In this letter, we study grand-canonical symmetric orbifolds of conformal field theories. We propose to define them as the direct sum of symmetric orbifolds of all degrees. The natural basis of operators is one that mixes all sectors. We describe this basis in terms of partial permutations, and explain how to define and calculate the operator product expansion in that framework. Our construction provides a conformal field theory interpretation of the central charge operator $\mathcal{I}$ in $AdS_{3}$ string theory.
@@ I Introduction
String theory on an $AdS_{3}\times S^{3}\times T^{4}$ background with one unit of NS-NS flux is dual to a grand-canonical ensemble of symmetric orbifold conformal field theories [1, 2, 3, 4, 5]. So far, these ensembles have been defined solely in terms of their correlation functions. Up to a normalization factor, these read
$\langle\cdot\rangle_{p}:=\sum_{d}p^{d}\langle\cdot\rangle^{\rm un}_{d}\,$, (1)
where $\langle\cdot\rangle^{\rm un}_{d}$ denotes the unnormalized correlation function in the symmetric orbifold of degree $d$. The purpose of this letter is to clarify the structure of these ensembles. We will define them as the direct sum of symmetric orbifolds of all degrees, thus enlarging their Hilbert space beyond that of a single symmetric orbifold in the large $N$ limit. This immediately implies that they are consistent, conformally invariant quantum field theories. The basis of operators of physical interest is one that mixes all sectors, and we will focus on spelling out its properties. In particular, we will explain how to take operator product expansions directly in this basis. We will be especially interested in the combinatorial properties of the operator that maps to the central charge operator $\mathcal{I}$ of the dual string theory.
@@ I.2 Road map
where $Z_{p}$ denotes the partition function of the grand-canonical theory. At each degree, one can take the operator product expansion of the twist operators. However, the corresponding OPE coefficients depend on the degree. For example, $C_{kk}^{\quad 0}(d)=\frac{d!}{k(d-k)!}$. Therefore, it is unclear whether an operator product expansion can be defined in the grand-canonical theory. To prove that such an expansion would be inconsistent, let us compute the grand-canonical two-point function explicitly on the sphere:
@@
An alternative way to describe the central charge operator $\mathcal{I}$ is as a differential operator with respect to the chemical potential [5]. As such, it generates an algebra of operators absent from the Hilbert space of the symmetric orbifold. Therefore, if the central charge operator of the string theory admits a field-theoretic dual, then its description requires the Hilbert space of the grand-canonical theory to be substantially enlarged. Following the intuition gained from the study of topological symmetric orbifolds [12, 13, 14], we will propose that the algebraic structure required for this enlargement is that of partial permutations [15]. A partial permutation is a permutation of a subset of $n$ elements. It is composed of a support $\mathcal{S}\subset\{1,\dots,n\}$ and a permutation $\rho$ of $\mathcal{S}$. We denote it as $(\mathcal{S},\rho)$, or alternatively, with the Latin letter $r$. The product of two partial permutations is obtained by multiplying the permutations and concatenating their supports:
@@
We define the orbit sum $A_{[r]}$ as the sum over all partial permutations in the conjugation orbit of $r$. These orbit sums are characterized by the size of their support and their cycle type. Our proposal is that in a grand-canonical theory, twist operators are indexed not by orbits of the infinite symmetric group, but rather by orbits of partial permutations. An immediate consequence of this proposal is that there are many central operators. In section III, we will identify these operators within the algebra generated by the dual of the central charge operator $\mathcal{I}$. In the next section, we focus on clarifying the structure of the grand-canonical theory. We define it as the infinite direct sum of symmetric orbifolds of all degrees, study the consequences of this definition, and identify the physical basis of states using the language of partial permutations.
@@ II.1 The Hilbert Space
To construct the Fock space of quantum field theory, one must sum over all $n$-particle Hilbert spaces. Similarly, the Hilbert space of the grand-canonical symmetric orbifold is obtained by summing over symmetric orbifold Hilbert spaces $\mathcal{H}^{(d)}$ of all degrees:
$\mathcal{H}=\bigoplus_{d\geq 0}\mathcal{H}^{(d)}\,.$ (15)
Each fixed-degree Hilbert space is itself the sum of twisted sectors, which are generated by the twist operators $\sigma_{[g]}^{(d)}$. In the grand-canonical context, we are interested in twists that act at each degree simultaneously, such as
$\alpha_{k}:=\bigoplus_{d\geq k}\sigma_{k}^{(d)}$ (16)
However, there are far too few operators of this form to constitute a basis of the full Hilbert space. To remedy this problem, we generalize the definition (16) to an arbitrary partial permutation:
$\alpha_{(\mathcal{S},\rho)}:=\bigoplus_{d\geq\lvert\mathcal{S}\rvert}\sigma_{\rho}^{(d)}\,$, (17)
The corresponding gauge-invariant twists then read
$\alpha_{r}:=\bigoplus_{d\geq\lvert d_{r}\rvert}\binom{m_{1}(\rho)}{m_{1}(r)}\sigma_{\rho}^{(d)}\,.$ (18)
where $m_{1}(r)$ is the number of trivial elements in the support of the partial permutation $r$, and $m_{1}(\rho)$ the number of trivial elements in the corresponding permutation $\rho$ of $\{1,2,\dots,d\}$. This combinatorial factor corresponds to the number of partial permutations $(\mathcal{S},\rho)$ that lift to $\rho$ in $S_{d}$ under a support-forgetting morphism [15]. The Hilbert space decomposes as
@@
where $\mathcal{H}_{[r]}$ is the sector of the Hilbert space whose ground state is generated by the twist $\alpha_{r}$, and the sum runs over orbits of partial permutations of all sizes. In this picture, each seed theory operator $X_{I}$ should be associated with a twist $\alpha_{(I)}$, which means we consider a copy of the seed operator at each degree $d>I$. Furthermore, the identity operator is the partial permutation orbit of empty support $\alpha_{\emptyset}$.
@@ II.2 Correlation Functions
We define the grand-canonical ensemble of symmetric orbifolds as the direct sum of symmetric orbifolds of all degrees. The one-point function is given by a weighted sum of the one-point function in each theory, weighted by a chemical potential $p$:
$\langle\cdot\rangle_{p}:=\sum_{d}p^{d}\langle\cdot\rangle^{\rm un}_{d}\,$, (20)
The recipe for computing correlation functions in the physical basis of operators follows immediately from the definition of partial permutations in terms of fixed-degree twists:
$\langle\alpha_{[r]}(z)\cdot(\dots)\rangle_{d}=\binom{m_{1}(\rho)}{m_{1}(r)}\langle\sigma_{[\rho]}(z)\cdot(\dots)\rangle_{d}\,.$ (21)
Geometrically, the operator $\alpha_{[r]}$ selects $\lvert\mathcal{S}_{r}\rvert$ sheets out of the $d$ available in the covering space. Only those sheets contribute to connected correlators. This geometric picture is perhaps best understood from the dual string perspective, where trivial elements in the support of a partial permutation correspond to disjoint worldsheets that cover the target space exactly once, see section III.4.
We note that combinatorial differences in fixed-degree correlation functions can lead to drastically different behaviours once summed over all degrees. For example, the operator $\alpha_{1}$ behaves within any fixed-degree correlation function as a factor of the degree:
@@ II.3 Operator Product Expansion
The definition and consistency of an associative operator product expansion in the grand-canonical theory follow immediately from its identification as a direct sum of conformal field theories. However, this alone does not guarantee that the coefficients appearing in the expansion of the product of two twist operators $\alpha_{r}$ are finite, since the change of basis involves infinite sums of states. Their finiteness instead follows from the finiteness of all grand-canonical correlation functions on the sphere. The latter is guaranteed by the fact that the chemical potential decreases exponentially with the degree $d$, whereas symmetric orbifolds correlators are only polynomial in $d$. We stress that these OPE coefficients are independent of the degree, since they were not defined with respect to any specific value of $d$.
Although the existence of the operator product expansion is guaranteed, it is only natural to ask for an intrinsic definition in the language of partial permutations. We now provide such a definition. Recall that to each seed theory field $X_{I}$, we associate a twist operator $\alpha_{(I)}$. Therefore, given the OPE coefficients of the seed theory, the problem reduces to computing products of twist operators $\alpha_{(\mathcal{S}_{1},\rho_{1})}$ and $\alpha_{(\mathcal{S}_{2},\rho_{2})}$, and regrouping the resulting terms. The product lies in a single channel $\alpha_{(\mathcal{S}_{1}\cup\mathcal{S}_{2},\sigma_{\rho_{1}}\sigma_{\rho_{2}})}$. Within this channel, the various contributions are obtained by embedding the twist operators into the infinite symmetric orbifold and computing the operator product expansion of $\sigma_{\rho_{1}}$ with $\sigma_{\rho_{2}}$. In fact, the computation may be carried out in a symmetric orbifold of any degree $d\geq\lvert\mathcal{S}_{1}\rvert+\lvert\mathcal{S}_{2}\rvert-1$. Indeed, its result is independent of spectator copies, that lie outside the support of the twists. Finally, one groups together appropriate contributions from each channel.
A simple algorithm consists of first computing the operator product expansion at some fixed degree, then replacing each factor of the degree by an operator $\alpha_{1}$, and finally multiplying out the factors of $\alpha_{1}$ using equations (37) and (40). Equivalently, polynomials of the degree may be decomposed in the basis of binomial polynomials $p_{k}(d):=\binom{d}{k}$, and each $p_{k}$ replaced with an operator $\alpha_{1^{k}}$. For instance, given a free scalar field as seed theory, the operator product expansion of the field $X:=\sum_{I}X_{I}$ with itself reads
$\partial X(z)\,\partial X(w)=$ $-\frac{d}{(z-w)^{2}}+\dots$
$\rightarrow$ $-\frac{\alpha_{1}}{(z-w)^{2}}+\dots\,$,
seeing as $\alpha_{1}$ behaves as a factor of the degree in any fixed-degree correlation function. In our framework, one associates a trivial twist $\alpha_{(I)}$ to each copy of the seed field $X_{I}$, and the resulting $\alpha_{1}$ arises from the sum over all $\alpha_{(I)}$. To conclude this section, we note that the partial permutation framework offers an illuminating perspective on operator product expansions in the large $N$ limit of symmetric orbifolds [16].
II.4 Virasoro Algebra
The stress-energy tensor of the grand-canonical theory reads
@@ II.4 Virasoro Algebra
where the $L_{n}$ are the Laurent modes of the stress-energy tensor, $L_{n}:=\int dz\,z^{n+1}T(z)$. An infinite-dimensional generalization of Schur’s lemma implies that the representation at hand is reducible, as it admits a non-scalar central element. This is consistent with the definition of the grand-canonical ensemble as a direct sum of conformal field theories with distinct central charges.
The appearance of a non-scalar central charge in the representation (26) is consistent with the findings of [17] for the spacetime Virasoro algebra of $AdS_{3}$. In the tensionless context, it leads to the identification of $\alpha_{1}$ as the counterpart of the central charge operator $\mathcal{I}$ of the dual string theory. The expectation value of $c_{seed}\,\alpha_{1}$ therefore computes the Brown-Henneaux central charge [18]. This allows us to identify the relation between the string coupling and the chemical potential. The three-dimensional Newton constant for an $AdS_{3}\times S^{3}\times T^{4}$ background at $k=1$ reads [19]
$G_{N}^{(3)}=\frac{G_{N}^{(10)}}{\text{Vol}(S^{3})\cdot\text{Vol}(\mathbb{T}^{4})}=\frac{8\pi^{6}g_{s}^{2}R^{8}}{2\pi^{2}R^{3}\cdot(2\pi)^{4}v_{4}R^{4}}\,$,
@@
At this stage, let us summarize our construction. We have shown that the grand-canonical ensemble of symmetric orbifolds can be defined as the direct sum of orbifolds of all degrees. This ensures that the resulting theory is an ordinary quantum field theory, with a well-defined and associative operator product expansion. It is also unitary, provided the seed theory is. Crucially, the basis of operators that admits a physical interpretation mixes all sectors. We have explained how to take operator product expansions in this basis, except for two formulas describing the combinatorics of the operator $\alpha_{1}$. We establish them in the next section.
@@ III.1 The Central Charge Operator
$\langle\alpha_{1}\cdot(\dots)\rangle^{\rm un}$ $:=\sum_{d\geq 1}d\,p^{d}\langle(\dots)\rangle^{\rm un}_{d}$ (29)
$=p\partial_{p}\langle(\dots)\rangle^{\rm un}\,.$ (30)
@@ III.2 The Universal Subalgebra
$\langle\alpha_{1^{k}}\cdot(\dots)\rangle^{\rm un}$ $:=\sum_{d\geq 1}\binom{d}{k}p^{d}\langle(\dots)\rangle^{\rm un}_{d}\,$, (32)
$=\frac{p^{k}\partial_{p}^{k}}{k!}\langle(\dots)\rangle^{\rm un}\,.$ (33)
@@ III.3 OPE with the Central Charge Operator
with $k\geq 2$. A geometric interpretation of equation (43) is that $\alpha_{1}$ either marks an existing sheet or creates a new one. Holographically, the latter corresponds to adding a disjoint worldsheet in its vacuum state. A key property of the Hilbert space (19) is indeed that it distinguishes between vacuum configurations with different numbers of worldsheets in their vacuum state. The string number is equal to the number of partial permutation orbits in the dual state. For instance, $\alpha_{1,3}=\alpha_{(1)(234)}+\text{etc.}$ represents a two-string state. Our construction suggests that for a boundary description to exhibit a field theory structure, and an operator product expansion in particular, it must holographically encode a multi-string Hilbert space.